1. Time Cash Flows 0 ($200,000) 1 0 2 0 3 0 4 0 5 0...

1.

Time

Cash Flows

0

($200,000)

1

0

2

0

3

0

4

0

5

0

6

0

7

0

8

0

9

0

10

$5,000,000

The internal rate of return of this cash flow stream is between 37% and 38%.

True

False

2.

Time

Cash Flows

0

($100,000)

1

0

2

0

3

0

4

0

5

0

6

0

7

0

8

0

9

0

10

0

11

0

12

$400,000

The internal rate of return of this cash flow stream is between 12% and 13%.

True

False

3.

Time

Cash Flows

0

($1,000,000)

1

0

2

0

3

0

4

0

5

0

6

0

7

0

8

0

9

0

10

$3,000,000

The internal rate of return of this cash flow stream is between 9% and 10%.

True

False

4.

Time

Cash Flows

0

($400,000)

1

0

2

0

3

0

4

0

5

0

6

0

7

0

8

0

9

0

10

$5,000,000

The internal rate of return of this cash flow stream is between 25% and 27%.

True

False

5.

Time

Cash Flows

0

($100,000)

1

0

2

0

3

0

4

400,000

The internal rate of return of this cash flow stream is between 41% and 42%.

True

False
Answer #1

Part 1:

FV = PV (1+r)^n

5000000 = 200000 (1+r)^10

(1+r)^10 = 5000000 / 200000

= 25

1+r = 25^(1/10)

= 1.3797

r = 1.3797 - 1.

= 0.3797 i.e 37.97%

True

Part 2:

FV = PV (1+r)^n

400000 = 100000 (1+r)^12

(1+r)^12 = 400000 / 100000

= 4

1+r = 4^(1/12)

= 1.1225

r = 1.1225 - 1.

= 0.1225 i.e 12.25%

True

Part 3:

FV = PV (1+r)^n

3000000 = 1000000 (1+r)^10

(1+r)^10 = 3000000 / 1000000

= 3

1+r = 3^(1/10)

= 1.1161

r = 1.1161 - 1.

= 0.1161 i.e 11.61%

False

Part 4:

FV = PV (1+r)^n

5000000 = 400000 (1+r)^10

(1+r)^10 = 5000000 / 400000

= 12.5

1+r = 12.5^(1/10)

= 12873

r = 1.2873 - 1.

= 0.2873 i.e 28.73%

False

Part 5:

FV = PV (1+r)^n

400000 = 100000 (1+r)^4

(1+r)^4 = 400000 / 100000

= 4

1+r = 4^(1/4)

= 1.4142

r = 1.4142 - 1.

= 0.4142 i.e 41.42%

True